f(x)=4x+2 and g(x)=3x-1 find (f-g)(4)
1
Expert's answer
2021-10-13T17:36:29-0400

As, (f-g(4))= f(4)-g(4)

As, f(x)= 4x+2

f(4)= 4(4)+2= 18

also,

g(x)= 3x-1 = 3(4)-1= 11


So, (f-g)(4)= f(4)-g(4)= 18-11


Hence ,

(f-g)(4) = 7